# Free-Lunch Color-Texture Disentanglement for Stylized Image Generation

### v2

@@ 4.3 Additional Applications
Figure 11: SADis allows for real-image based stylization (left). Besides, SADis effectively disentangles color and material elements from separate images, enabling precise control over color and material in image generation.
@@
Color and material transfer. Given color and material references from separate images, SADis is also capable of disentangling color and material elements, and provides precise controls over color and materials in T2I generation. The color and material transfer results are shown in Fig. 11 (right). This capability of SADis highlights its potential applications in artistic creation and industrial design.
@@ 5 Conclusion
In this paper, we address a critical challenge in stylized image generation by introducing the concept of disentangled stylized image generation (DisIG). We focus on color and texture as key style attributes, presenting the first approach, named style attribute disentanglement (SADis), for independent control over these elements. Using the image-prompt additivity property, we proposed novel techniques, including the color-texture extraction (CTE) and regularized whitening-coloring transformation (RegWCT), to ensure enhanced color-texture consistency and more accurate results. Experimental evaluations demonstrate that SADis significantly outperforms existing stylization methods, both qualitatively and quantitatively. This work not only advances the field of T2I stylization but also opens new avenues for more flexible and customizable image generation, paving the way for future innovations for art creators.
@@ Acknowledgements
We acknowledge project PID2022-143257NB-I00, financed by MCIN/AEI/10.13039/501100011033 and FEDER, and the Generalitat de Catalunya CERCA Program. This work was also supported by NSFC (NO. 62225604) and Youth Foundation (62202243). We acknowledge “Science and Technology Yongjiang 2035” key technology breakthrough plan project (2024Z120).
Supplementary Material: Free-Lunch Color-Texture Disentanglement for Stylized Image Generation
Appendix S1 Statements
@@
Appendix S2 Image-Prompt Additivity
Figure S1: By subtracting or adding the object image embeddings from the input image embeddings, we can effectively remove or add the object to the scene, although some degree of identity information for the person is also diminished.
Figure S2: (Left) By subtracting the grayscale image embeddings from the color image embeddings, we can effectively remove the texture information and isolate the color information. Combined with the prompt “A photo of a bird” we can generate images in the same color schemes as the input images. (Right) By subtracting the bluish image embeddings with the grayscale embeddings for random generated images with 100 prompts, we visualize the subtracted embedding via PCA decomposition. We observe them gathering around the pure blue image. The generation with these subtracted embeddings further prove the consistency.
Figure S3: By subtracting the color image embeddings with the grayscale embeddings for random generated images with 100 prompts, we visualize the subtracted embedding via PCA decomposition. We observe them gathering around the color reference image. The generation with these subtracted embeddings further prove the consistency.
S2.1 Broader Property of Image-Prompt Additivity
We demonstrate the broader property of Image-Prompt Additivity in Fig. S1. For instance, in Fig. S1-(a), subtracting the embedding of a hat from the embedding of a person wearing a hat results in the generation of a person without a hat. Similarly, in Fig. S1-(b), adding the embedding of glasses to a person results in the generation of that person wearing glasses. We refer to this phenomenon as Image-Prompt Additivity. We hypothesize this property originates from the image-text paired training in CLIP models [55]. The training process endows the image branch with the additivity property inherent to the text embedding space [51, 33, 6]. Although the person identities are altered after embedding additivity manipulations, this approach demonstrates a promising property for enabling our training-free color-texture disentanglement. Building on this property, we develop our method, SADis, to extract color and texture information from reference images and effectively apply these attributes in T2I generation.
@@
S2.2 Extra Analysis on Image-Prompt Additivity
@@
As another illustration of image prompt additivity, we construct a set of blue, red, and green images. For example, the blue set is constructed by setting the red and green color channels to zero of a set of 100 images (similarly, for the green and red set). We isolate the color representation by performing a feature subtraction between the blue image and its grayscale equivalent. We plot their projected embeddings as blue dots in Fig. S2-(Right); note how these embeddings always maintain close to the pure blue embedding and are far from the green and red embeddings. Also, generated images with these subtraction blue color embeddings keep the bluish color palette as expected (see the car and t-shirt). An extended analysis is presented in Fig. S3, where we apply similar manipulations using colorful images as color references. The successful clustering further confirms that the resulting embeddings effectively preserve essential color information.
Appendix S3 Implementation details
@@
We develop our method, SADis, based on the SDXL model [53], which is among the leading open-source T2I generative models available. To construct the experimental datasets, we randomly select 40 images from the WikiArt dataset [64] and StyleDrop [62] image collections. Each image can be used either as a color reference or a texture reference in the experiments. All input images are resized to $512\times 512$ before feeding into models. Also, for quantitative comparisons, 20 images are designated as the color reference set, while the remaining 20 serve as the texture reference set. Additionally, we sample 10 content prompts from StyleDrop, yielding a total of 4000 stylized images for each method for comparison. Since some comparison methods (such as SDXL [53] and Artist [39]) require text prompts as style controls for image generation, we employ the state-of-the-art vision-language model GPT-4o as the image captioning tool to generate precise color and texture prompts from the reference images. Subsequently, the content, color, and texture prompts are concatenated in the format: {content prompt}, {texture prompt} in the color {color prompt}. As for the comparison models (like DEADiff [54], IP-Adapter [75], Instantstyle [67], DreamStyler [3], StyleDrop [62], and CSGO [73]) that require text prompts for color control, the text prompt is constructed as: {content prompt}, in the color {color prompt}.
Figure S4: Some color and texture prompts generated by GPT-4o for comparison methods in the main paper for qualitative evaluation.
@@
We further include the details of each evaluation metric. CLIP Score [55] is used to evaluate the semantic alignment between the text prompt and the generated image. We calucate the CLIP Score according to configuration of the T2I-CompBench repository [34]. For the color alignment evaluation, we use MS-SWD [25], color histogram distance (C-Hist), and GPT4o to calculate the color attribute similarity between the color reference and the generated image. MS-SWD [25]: since the generated imgae are usually not spatially aligned with the color reference image, we use MS-SWD to better evaluate the color attribute alignment according to the default setting of their repository [25]. C-Hist: we first compute the RGB histograms of the color reference and generated images. Afterwards, the Bhattacharyya distance is used to measure the differences between their color histograms. GPT-4o score [1]: to comprehensively evaluate the color alignment performance, we furture adopt the multimodal model GPT-4o to compute the color alignment score between the generated image and the color reference image. Specifically, the GPT-4o metric is computed according to Fig. S5. Firstly, following the previous work[17], we extract dominant colors by ColorThief. Afterwards, we feed the extracted color names to GPT-4o as the reference color from the color image, and ask GPT-4o to give 1-5 points according to the criteria shown in Fig. S5.
Figure S5: GPT-4o metric for color alignment evaluation.
Appendix S4 WCT transformations formulas
@@
Given a latent $z_{t}^{{}^{\prime}}\in\mathcal{R}^{C\times H\times W}$ and a color reference latent $z_{t}^{c}\in\mathcal{R}^{C\times H\times W}$ at timestep $t$, we adopt Whitening and Coloring Transforms (WCT) [44, 10, 31] to transform $z_{t}^{{}^{\prime}}$ to match covariance matrix of color reference latent $z_{t}^{c}$. There are two step for WCT: Whitening transform and Coloring transform.
@@
Coloring transform: it’s the reverse process of Whitening transform [32, 44]. Beforehand, the color reference latent $z_{t}^{c}$ is centered by subtracting its mean vector $m_{c}$. Subsequently, we obtain the transformed latent $z_{t}^{wct}$ which satisfies the desired correlations $z_{t}^{wct}{z_{t}^{wct}}^{\text{T}}=z_{t}^{c}{z_{t}^{c}}^{\text{T}}=I$:
@@
Appendix S5 Additional Experimental Results
@@
In Fig. S6, we include more color-texture disentanglement examples of our method SADis, which is under the Disentangled Stylized Image Generation (DisIG) scenario.
Figure S6: Additional experimental results of SADis.
@@
Image-based stylization. Furthermore, our method, SADis, can be seamlessly integrated with ControlNet [77] to enable image-based stylized generation, as demonstrated in Fig. S7. Here, we adopt ControlNet (Canny) as the base pipeline, setting its conditioning scale to 0.6. All other hyper-parameters are kept consistent with those of SADis. This integration broadens the application scenarios of our proposed approach.
@@
Figure S7: Our work is compatible with ControlNet to achieve image-based stylization.
@@
Color and material transfer. Given color and material references from separate images, our method, SADis, is also capable of disentangling color and material elements, and provides precise controls over color and materials in T2I generation. The color and material transfer results are shown in Fig. S9. This capability of SADis highlights its potential applications in artistic creation and industrial design. We also desire to note that, the term “color palette” typically refers to the overall color scheme of an image, not specific local regions. Our SADis does not control the color palette of individual objects in general. However, Fig. S9 demonstrates controlled color and material transfer using a masking mechanism with our method SADis.
@@
Figure S8: SADis can control color and texture generation separately using reference images.
@@
Figure S9: SADis effectively disentangles color and material elements from separate images, enabling precise control over color and material in image generation.
@@
Robust to color variations. As shown in Fig.S10 and Fig. S11, we vary the saturation and illuminant continuously. To be specific, for saturation, we developed a Python-based saturation adjustment tool utilizing HSV (Hue, Saturation, Value) color space transformation. The tool linearly modifies the saturation channel using factors of [0.2, 0.6, 1.0, 1.5, 3.0], respectively, before converting the images back to RGB format. The results of saturation variations are shown in Fig. S10. Regarding the color of the illuminant, we modify images by applying calibrated RGB channel multipliers. For warm temperatures (red) as shown in Fig. S11, red increases by 30%, green reduces by 10%, and blue by 20%. The intensity of these adjustments can be controlled (0-1 range), enabling fine-grained color temperature manipulation. We present 3 increasing intensities in Fig. S11. Our method SADis is able to generate images with smooth change along with the saturation and illuminant color variations, while the other method InstantStyle [67] fails. That further proves the robustness of our method SADis for Disentangled Stylized Image Generation (DisIG).
@@
Figure S10: Compared to InstantStyle, SADis preserves the content more effectively under saturation variations.
@@
Figure S11: Compared to InstantStyle, SADis preserves the content more effectively under illuminant variations.
@@
Each component works independently. The T2I generations in Fig. S8-left provide evidence that the color and texture branches in SADis function independently. Additionally, the visualization for our ablatio study in  Fig. S8 further demonstrates improved color alignment with minimal texture degradation, achieving a good trade-off. However, the trade-off between texture and color alignment is optional for users, allowing them to adjust the balance based on the requirements of specific application scenarios.
Appendix S6 Ablation studies on RegWCT

### v3

@@ 4.3 Additional Features
Figure 7: (a) SADis allows for real-image stylization. (b) SADis effectively disentangles color and material elements from separate images, enabling precise control over color and material in image generation. (c) SADis remains effective when exchanging texture and color images. (d) SADis can capture color image properties such as saturation and illuminant. This further demonstrates the robustness of SADis in realistic DisIG scenario.
Image-based stylization. SADis Integrated with ControlNet achieves image-based stylized generation, as shown in Fig.˜7 (a). Here, we adopt ControlNet (Canny) as the base model, setting its conditioning scale to 0.6. This integration broadens the application scenarios of SADis.
@@
Color and material transfer. Given color and material references from separate images, SADis is also capable of disentangling color and material elements, and provides precise controls over color and materials in T2I generation. The color and material transfer results are shown in Fig.˜7 (b). This capability of SADis highlights its potential applications in artistic creation and industrial design.
Compatible with MMDiT-based models. SADis is also compatible with MMDiT-based models (SD3.5 sd3.5 and FLUX.1-dev [^5] https://huggingface.co/black-forest-labs/FLUX.1-dev.) to achieve color-texture disentangled stylization. Specifically, we use SD3.5 and FLUX.1-dev as the base models. These models employ the original text encoders (i.e., T5, CLIP-L, and CLIP-G) for content control. We utilize SigLIP siglip with a pretrained adapter [^6] https://huggingface.co/InstantX/SD3.5-Large-IP-Adapter., [^7] https://huggingface.co/InstantX/FLUX.1-dev-IP-Adapter. for style control. Color-texture embeddings are extracted and disentangled from SigLIP features, following the same workflow as illustrated in Fig.˜3. As shown in Fig.˜8, SADis achieves color-texture disentanglement and enables stylized image generation on these MMDiT-based models as well. This demonstrates that the color-texture disentanglement capability of SADis does not rely on the U-Net architecture and can be generalized to MMDiT-based models for stylized image generation. It also indicates that SADis is not limited to specific CLIP models and can be applied to other text-image pretrained encoders.
Figure 8: SADis is compatible with MMDiT-based models. (a) demonstrates color-texture disentangled stylization results on FLUX.1-dev, while (b) shows the corresponding results on SD3.5.
@@ 5 Conclusion
We addressed a critical challenge in stylized image generation by introducing the concept of disentangled stylized image generation (DisIG). We focused on color and texture as key style attributes, presenting the first approach, named style attribute disentanglement (SADis), for independent control over these elements. Using the image-prompt additivity property, we proposed novel techniques, including the color-texture extraction (CTE) and regularized whitening-coloring transformation (RegWCT), to ensure enhanced color-texture consistency and more accurate results. Experimental evaluations demonstrate that SADis significantly outperforms existing stylization methods, both qualitatively and quantitatively. This work also opens new avenues for more flexible and customizable image generation, paving the way for future innovations for art creators.
@@ Acknowledgements
We acknowledge project PID2022-143257NB-I00, financed by MCIN/AEI/10.13039/501100011033 and ERDF/EU, and the Generalitat de Catalunya CERCA Program, and ELLIOT project funded by the European Union under Grant Agreement 101214398. This work was also supported by NSFC (NO. 62225604) and Youth Foundation (62202243). We acknowledge “Science and Technology Yongjiang 2035” key technology breakthrough plan project and Chinese government-guided local science and technology development fund projects (scientific and technological achievement transfer and transformation projects) (254Z0102G). Kai Wang acknowledges the funding from Guangdong and Hong Kong Universities 1+1+1 Joint Research Collaboration Scheme and the start-up grant B01040000108 from CityU-DG.
Appendix
Appendix A Statements
@@
Appendix B Image-Prompt Additivity
B.1 Broader Property of Image-Prompt Additivity
We demonstrate the broader property of Image-Prompt Additivity in Fig.˜9. For instance, in Fig.˜9-(a), subtracting the embedding of a hat from the embedding of a person wearing a hat results in the generation of a person without a hat. Similarly, in Fig.˜9-(b), adding the embedding of glasses to a person results in the generation of that person wearing glasses. We refer to this phenomenon as Image-Prompt Additivity. We hypothesize this property originates from the image-text paired training in CLIP models radford2021clip . The training process endows the image branch with the additivity property inherent to the text embedding space NIPS2013_wordembedding ; hu2024tome ; manuel2023sega . Although the person identities are altered after embedding additivity manipulations, this approach demonstrates a promising property for enabling our training-free color-texture disentanglement. Building on this property, we develop our method, SADis, to extract color and texture information from reference images and effectively apply these attributes in T2I generation.
@@ B.1 Broader Property of Image-Prompt Additivity
Figure 9: By subtracting or adding the object image embeddings from the input image embeddings, we can effectively remove or add the object to the scene, although some degree of identity information for the person is also diminished.
Figure 10: (Left) By subtracting the grayscale image embeddings from the color image embeddings, we can effectively remove the texture information and isolate the color information. Combined with the prompt “A photo of a bird” we can generate images in the same color schemes as the input images. (Right) By subtracting the bluish image embeddings with the grayscale embeddings for random generated images with 100 prompts, we visualize the subtracted embedding via PCA decomposition. We observe them gathering around the pure blue image. The generation with these subtracted embeddings further prove the consistency.
Figure 11: By subtracting the color image embeddings with the grayscale embeddings for random generated images with 100 prompts, we visualize the subtracted embedding via PCA decomposition. We observe them gathering around the color reference image. The generation with these subtracted embeddings further prove the consistency.
B.2 Extra Analysis on Image-Prompt Additivity
@@
As another illustration of image prompt additivity, we construct a set of blue, red, and green images. For example, the blue set is constructed by setting the red and green color channels to zero of a set of 100 images (similarly, for the green and red set). We isolate the color representation by performing a feature subtraction between the blue image and its grayscale equivalent. We plot their projected embeddings as blue dots in Fig.˜10-(Right); note how these embeddings always maintain close to the pure blue embedding and are far from the green and red embeddings. Also, generated images with these subtraction blue color embeddings keep the bluish color palette as expected (see the car and t-shirt). An extended analysis is presented in Fig.˜11, where we apply similar manipulations using colorful images as color references. The successful clustering further confirms that the resulting embeddings effectively preserve essential color information.
Appendix C Implementation details
@@
We develop our method, SADis, based on the SDXL model podell2023sdxl , which is among the leading open-source T2I generative models available. To construct the experimental datasets, we randomly select 40 images from the WikiArt dataset artgan2018_wikiart and StyleDrop sohn2023styledrop image collections. Each image can be used either as a color reference or a texture reference in the experiments. All input images are resized to $512\times 512$ before feeding into models. Also, for quantitative comparisons, 20 images are designated as the color reference set, while the remaining 20 serve as the texture reference set. Additionally, we sample 10 content prompts from StyleDrop, yielding a total of 4000 stylized images for each method for comparison. Since some comparison methods (such as SDXL podell2023sdxl and Artist jiang2024artist ) require text prompts as style controls for image generation, we employ the state-of-the-art vision-language model GPT-4o as the image captioning tool to generate precise color and texture prompts from the reference images. Subsequently, the content, color, and texture prompts are concatenated in the format: {content prompt}, {texture prompt} in the color {color prompt}. As for the comparison models (like DEADiff qi2024deadiff , IP-Adapter ye2023ip-adapter , Instantstyle wang2024instantstyle , DreamStyler ahn2023dreamstyler , StyleDrop sohn2023styledrop , and CSGO xing2024csgo ) that require text prompts for color control, the text prompt is constructed as: {content prompt}, in the color {color prompt}. All experiments were conducted on an NVIDIA-L40s GPU.
@@
We further include the details of each evaluation metric. CLIP Score radford2021clip is used to evaluate the semantic alignment between the text prompt and the generated image. We calucate the CLIP Score according to configuration of the T2I-CompBench repository t2i-benchmark . For the color alignment evaluation, we use MS-SWD he2024_MS_SWD , color histogram distance (C-Hist), and GPT4o to calculate the color attribute similarity between the color reference and the generated image. MS-SWD he2024_MS_SWD : since the generated imgae are usually not spatially aligned with the color reference image, we use MS-SWD to better evaluate the color attribute alignment according to the default setting of their repository he2024_MS_SWD . C-Hist: we first compute the RGB histograms of the color reference and generated images. Afterwards, the Bhattacharyya distance is used to measure the differences between their color histograms. GPT-4o score achiam2023chatgpt4 : to comprehensively evaluate the color alignment performance, we furture adopt the multimodal model GPT-4o to compute the color alignment score between the generated image and the color reference image. Specifically, the GPT-4o metric is computed according to Fig.˜12. Firstly, following the previous workgal2022textual , we extract dominant colors by ColorThief. Afterwards, we feed the extracted color names to GPT-4o as the reference color from the color image, and ask GPT-4o to give 1-5 points according to the criteria shown in Fig.˜12.
Figure 12: GPT-4o metric for color alignment evaluation.
Figure 13: Some color and texture prompts generated by GPT-4o for comparison methods in the main paper for qualitative evaluation.
Appendix D WCT transformations formulas
@@
Given a latent $z_{t}^{{}^{\prime}}\in\mathcal{R}^{C\times H\times W}$ and a color reference latent $z_{t}^{c}\in\mathcal{R}^{C\times H\times W}$ at timestep $t$, we adopt Whitening and Coloring Transforms (WCT) li2017universal ; cho2019i2i_wct ; hossain2014wct to transform $z_{t}^{{}^{\prime}}$ to match covariance matrix of color reference latent $z_{t}^{c}$. There are two step for WCT: Whitening transform and Coloring transform.
@@
Coloring transform: it’s the reverse process of Whitening transform wct ; li2017universal . Beforehand, the color reference latent $z_{t}^{c}$ is centered by subtracting its mean vector $m_{c}$. Subsequently, we obtain the transformed latent $z_{t}^{wct}$ which satisfies the desired correlations $z_{t}^{wct}{z_{t}^{wct}}^{\text{T}}=z_{t}^{c}{z_{t}^{c}}^{\text{T}}=I$:
@@
Appendix E Additional Experimental Results
@@
In Fig.˜14, we include more color-texture disentanglement examples of our method SADis, which is under the Disentangled Stylized Image Generation (DisIG) scenario.
Figure 14: Additional experimental results of SADis.
Figure 15: Stylized images generated by our training-free method, SADis. (Up) As shown in the first rows, it enables disentangled control over color and style attributes in text-to-image diffusion models using separate image prompts. This approach offers creators enhanced color control, including the use of color palettes as in the last two columns. (Down) SADis also enables real-image stylization by incorporating a content image as an additional condition via ControlNet. Furthermore, it extends to color-only stylized generation and material transfer for more flexible image generation.
Additional comparison results with complex prompts. To further evaluate performance with complex prompts, we conducted additional experiments by randomly sampling 10 complex prompts from DREAMBENCH++ peng2024dreambench , together with 10 color and 10 texture images from our dataset, resulting in 1000 generated images for evaluation. As shown in Tab.˜3, compared to other methods, our approach achieves the best disentanglement of color and texture, resulting in a more balanced performance in both color and texture consistency. Specifically, our method attains significantly higher color consistency, while other methods often exhibit strong color-texture entanglement, with color being overly influenced by the texture reference image. Although IP-Adapter achieves the highest numerical score for texture consistency, it suffers from severe semantic leakage from the texture reference image (Row 2 and Row 3 in Fig.˜6), leading to poor text-image alignment (CLIP score: 0.257), which is inadequate for stylization. In contrast, except for IP-Adapter, our method achieves the highest scores in both texture consistency and text-image alignment.
Table 3: Quantitative comparison with complex prompts. The complex prompts are sampled from DREAMBENCH++ peng2024dreambench . The best and second-best numbers are marked with bold and underlined respectively.
Method
CLIP
Color Texture
[…]
SDXL podell2023sdxl 0.291 9.00 1.14 3.07 0.698 0.091
IPAdapter ye2023ip-adapter 0.257 9.99 1.15 2.96 0.817 0.058
InstantStyle wang2024instantstyle 0.278 11.21 1.25 2.85 0.751 0.065
Artist jiang2024artist 0.253 11.15 1.24 2.67 0.742 0.082
DEADiff qi2024deadiff 0.280 10.01 1.14 2.80 0.718 0.090
StyleDrop sohn2023styledrop 0.290 11.90 1.35 2.83 0.731 0.078
DreamStyler ahn2023dreamstyler 0.300 11.27 1.21 2.57 0.677 0.083
CSGO xing2024csgo 0.299 12.60 1.22 2.64 0.728 0.081
SADis (Ours)
0.301 6.08 0.95 3.14 0.751 0.064
@@
Image-based stylization. Furthermore, our method, SADis, can be seamlessly integrated with ControlNet zhang2023controlnet to enable image-based stylized generation and material transfer, as demonstrated in Fig.˜16 and Fig.˜15(the last row) respectively. Here, we adopt ControlNet (Canny) as the base pipeline, setting its conditioning scale to 0.6. All other hyper-parameters are kept consistent with those of SADis. This integration broadens the application scenarios of our proposed approach.
@@
Figure 16: Our work is compatible with ControlNet to achieve image-based stylization.
@@
Color and material transfer. Given color and material references from separate images, our method, SADis, is also capable of disentangling color and material elements, and provides precise controls over color and materials in T2I generation. The color and material transfer results are shown in Fig.˜19 and Fig.˜15(the last row). This capability of SADis highlights its potential applications in artistic creation and industrial design. We also desire to note that, the term “color palette” typically refers to the overall color scheme of an image, not specific local regions. Our SADis does not control the color palette of individual objects in general. However, Fig.˜19 demonstrates controlled color and material transfer using a masking mechanism with our method SADis.
Figure 17: Visualization of ablating each component of SADis.
@@
Figure 18: SADis can control color and texture generation separately using reference images.
@@
Figure 19: SADis effectively disentangles color and material elements from separate images, enabling precise control over color and material in image generation.
Color and texture transfer from the same image. SADis performs exceptionally well when using the same image as both the color and texture reference (as shown in Fig.˜20), showcasing its remarkable flexibility and adaptability. This ability highlights the method’s robustness in leveraging a single source to effectively guide both color and texture information, ensuring consistent and coherent results. The comparative experimental results are presented in Tab.˜4. In terms of texture consistency, our method achieves the best performance among all methods except for IP-Adapter. However, it is important to note that although the IP-Adapter achieves the highest numerical score for texture consistency, it introduces severe semantic leakage from the texture reference image (see Row. 2 and Row. 3 of IP-adapter results in Fig.˜6), resulting in poor text-image alignment (CLIP: 0.260 in Tab.˜4), which fails to meet the requirements of stylization. In contrast, excluding the IP-Adapter, our method achieves the highest scores in color consistency (MS-SWD: 3.19), texture consistency (CLIP-I: 0.754), and text-image alignment (CLIP: 0.302).
Table 4: Quantitative comparison where both color and texture are derived from the same image. The best and second-best numbers are marked with bold and underlined, respectively.
Method
CLIP
Color Texture
[…]
SDXL podell2023sdxl 0.292 9.12 1.15 3.04 0.696 0.090
IPAdapter ye2023ip-adapter 0.260 4.18 0.63 3.44 0.815 0.057
InstantStyle wang2024instantstyle 0.277 4.21 0.64 3.41 0.753 0.066
Artist jiang2024artist 0.272 10.98 1.18 2.83 0.744 0.077
DEADiff qi2024deadiff 0.295 10.81 1.14 2.66 0.721 0.089
StyleDrop sohn2023styledrop 0.292 12.16 1.28 2.51 0.717 0.090
DreamStyler ahn2023dreamstyler 0.301 11.90 1.09 2.53 0.691 0.095
CSGO xing2024csgo 0.298 6.43 0.83 3.09 0.716 0.089
SADis (Ours)
0.302 3.19 0.53 3.51 0.754 0.066
@@
Robust to color variations. As shown in Fig.21 and Fig.˜22, we vary the saturation and illuminant continuously. To be specific, for saturation, we developed a Python-based saturation adjustment tool utilizing HSV (Hue, Saturation, Value) color space transformation. The tool linearly modifies the saturation channel using factors of [0.2, 0.6, 1.0, 1.5, 3.0], respectively, before converting the images back to RGB format. The results of saturation variations are shown in Fig.˜21. Regarding the color of the illuminant, we modify images by applying calibrated RGB channel multipliers. For warm temperatures (red) as shown in Fig.˜22, red increases by 30%, green reduces by 10%, and blue by 20%. The intensity of these adjustments can be controlled (0-1 range), enabling fine-grained color temperature manipulation. We present 3 increasing intensities in Fig.˜22. Our method SADis is able to generate images with smooth change along with the saturation and illuminant color variations, while the other method InstantStyle wang2024instantstyle fails. That further proves the robustness of our method SADis for Disentangled Stylized Image Generation (DisIG).
@@
Figure 21: Compared to InstantStyle, SADis preserves the content more effectively under saturation variations.
@@
Figure 22: Compared to InstantStyle, SADis preserves the content more effectively under diverse illuminant variations.
@@
Each component works independently. The T2I generations in Fig.˜18-left provide evidence that the color and texture branches in SADis function independently. Additionally, the visualization for our ablation study in  Fig.˜17 further demonstrates improved color alignment with minimal texture degradation, achieving a good trade-off. However, the trade-off between texture and color alignment is optional for users, allowing them to adjust the balance based on the requirements of specific application scenarios.
Stability of different sampling rounds. We conducted additional experiments to assess the stability of color and texture in repeated generations. For each setting, the process was repeated five times, generating 1,000 images per round and computing the relevant metrics. As summarized in Tab.˜5, the results show minimal variation in color and texture consistency across rounds, demonstrating the robustness and stability of our method.
Table 5: Quantitative results for different sampling rounds.
Round
CLIP
Color Texture
[…]
1 0.301 6.08 0.95 3.14 0.751 0.064
2 0.301 6.05 0.95 3.15 0.750 0.065
3 0.300 6.23 0.96 3.12 0.752 0.064
4 0.301 6.16 0.95 3.12 0.751 0.066
5 0.304 6.08 0.95 3.14 0.749 0.066
Appendix F Ablation studies on RegWCT